# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

### v3

@@ III-A Optimal Control Problem
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (10)
[…]
$\;\;\mu_{0}=\overline{\mu}_{0}\;\;\mu_{T}=\mu_{f}$ (11)
@@ III-B Convex Relaxation
To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations. We first introduce the following change of variables to convexify eq. 7:
$L_{k}(i)$ $=K_{k}(i){\Sigma}_{k},\;$ $Y_{k}(i)$ $=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$ (12)
@@
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)L^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B^{\top}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}\right.$
$\left.+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)]\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$
For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a finite sequence of $(2N+1)$-tuples $\{\Sigma_{k},L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy section III-B. Thus, section III-B can be expressed as
@@
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}\succeq 0$ (16)
We use these new variables in the objective function as follows:
@@
$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right]$ (18a)
$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\Omega}\rho_{k}(i)\left[\text{Tr}({\Sigma}_{k}Q_{k}+Y_{k}(i)R_{k})\right]$ (18b)
@@ III-C Two-step Convex Optimization Framework
We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 4 and then uses its solution to solve Problem 5.
@@
$\Sigma_{0}=\overline{\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (20c)
@@
Under Assumptions 1, 2 and 3, the optimal solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
@@
$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (21)
@@
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (22a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (22b)
@@
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq. 22c. Rearranging eq. 22a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined earlier in eq. 12, we simplify eq. 22b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Taking the determinant of both sides, we get the following:
$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
@@ IV Chance-Constrained MJLS Covariance Steering
$P(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (23a)
$P(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (23b)
@@
$P(p^{\top}_{j}u_{k}(i)+q_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (24a)
$P(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (24b)
Previous works [24, 8] have focused on deterministic and convex reformulations of eqs. 23 and 24, but they usually require Gaussian dynamics. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution. To achieve this, we seek to minimize the Kullback-Leibler divergence [25] between the MJLS distribution and the approximating Gaussian. It has been shown previously [26] that this is achieved by performing moment matching, i.e. setting the first two moments of the approximating Gaussian to be the same as the original distribution which is being approximated. We leverage this useful property to approximate the state and control of MJLS using Gaussian distributions. We approximate $x_{k}$ by $\mathcal{N}(\mu_{k},\Sigma_{k})$, using moment matching to the approximating Gaussian. Similarly, for control in mode $i$, $\mathbb{E}[u_{k}(i)]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}]=K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)=Y_{k}(i)$, therefore $u_{k}(i)$ can be approximated by $\mathcal{N}(\overline{u}_{k}(i),Y_{k}(i))$. We now discuss the deterministic formulation of eqs. 23 and 24.
@@ IV-A Deterministic Formulation
Lemma 2 (Lemma 2 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$ and $\epsilon\in(0,0.5)$. Then, $\mathbb{P}[a_{j}^{\top}{\xi}+b_{j}\leq 0,\;\forall j]\geq 1-\epsilon$ is implied by
$a_{j}^{\top}\bar{{\xi}}+b_{j}+m_{\mathcal{N}}(\epsilon_{j})\sqrt{a_{j}^{\top}P_{\xi}a_{j}}\leq 0,\;\forall j,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (25)
where $m_{\mathcal{N}}(\epsilon)=Q_{X\sim\mathcal{N}}(1-\epsilon)$ denotes the quantile function of the standard normal distribution at probability $1-\epsilon$.
Lemma 3 (Lemma 3 in [8]).
Let ${\xi}\sim\mathcal{N}(\bar{{\xi}},P_{\xi})\in\mathbb{R}^{n_{\xi}}$, $\epsilon\in(0,1)$, and $\gamma>0$. Then, $\mathbb{P}[\|{\xi}\|\leq\gamma]\geq 1-\epsilon$ is implied by
$\|\bar{{\xi}}\|+m_{\chi^{2}}(\epsilon,n_{\xi})\sqrt{\lambda_{\max}(P_{\xi})}\leq\gamma$ (26)
$m_{\chi^{2}}(\epsilon,n_{\xi})=\sqrt{Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)}$. Here, $Q_{X\sim\chi^{2}(n_{\xi})}(1-\epsilon)$ denotes the quantile function of the chi-squared distribution with $n_{\xi}$ degrees of freedom, evaluated at probability $(1-\epsilon)$.
Note that while eqs. 25 and 26 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
$m^{2}_{\mathcal{N}}(\epsilon_{j}){a^{\top}_{j}P_{\xi}a_{j}}\leq(a_{j}^{\top}\bar{\xi}+b_{j})^{2}$ (27)
$m^{2}_{\chi^{2}}(\epsilon,n_{\xi}){\lambda_{\max}(P_{\xi})}\leq(\gamma-\|\bar{\xi}\|)^{2}$ (28)
which are convex in the covariance $P_{\xi}$.

### v4

@@ III-A Optimal Control Problem
Solve Problem 2 without constraints in eq. 2d.
@@ III-B Convex Relaxation
The nonlinear terms in eq. 7 result in a non-convex optimization landscape for Problem 3. We first relax the nonlinear terms involving the covariance variables $K_{k}(i)$ and $S_{k}(i)$ in eq. 7 by introducing the following change of variables:
$L_{k}(i)$ $=K_{k}(i){{S}_{k}(i)},\;$ $Y_{k}(i)$ $=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$ (10)
@@
$S_{k+1}(j)=\Bigl(\sum_{i\in\mathcal{S}}\left[p_{ij}\{A_{k}(i)S_{k}(i)A^{\top}_{k}(i)\right$.
$+A_{k}(i)L_{k}^{\top}(i)B_{k}^{\top}(i)+B_{k}(i)L_{k}(i)A_{k}^{\top}(i)+B_{k}(i)Y_{k}(i)B_{k}^{\top}(i)$
$+\rho_{k}(i)B_{k}(i)\overline{u}_{k}(i)(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
$+\rho_{k}(i)c_{k}(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
$+\rho_{k}(i)A_{k}(i)\overline{x}_{k}(i)(A_{k}(i)\overline{x}_{k}(i)+B_{k}(i)\overline{u}_{k}(i)+c_{k})^{\top}$
$+G_{k}(i)G_{k}^{\top}(i))\}\Bigr)-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (11)
For a given mean trajectory $\tau$, we define covariance trajectory $\Xi$ as a set of two finite sequences of $N+1$ and $2N$-tuples $\{\Sigma_{k}\}_{k\in\mathbb{Z}_{0:T}}$ and $\{L_{k}(1)\ldots L_{k}(N),Y_{k}(1)\ldots Y_{k}(N)\}_{k\in\mathbb{Z}_{0:T-1}}$ which satisfy section III-B. Thus, section III-B can be expressed as
@@
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{{S}_{k}(i)}\end{bmatrix}\succeq 0$ (14)
We note that section III-B still has nonlinear terms involving the mean variables. To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization scheme which solves the mean and covariance problem sequentially. The objective function can be divided into a mean and covariance part using the new variables as follows:
@@
$J(\tau)=\sum_{k=0}^{T-1}$ $\sum_{i\in\mathcal{S}}\rho_{k}(i)\{\mathrm{Tr}(\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)Q_{k}+\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\}$ (16a)
$J(\Xi)$ $=\sum_{k=0}^{T-1}\sum_{i\in\mathcal{S}}\left[\text{Tr}({S}_{k}(i)Q_{k}+Y_{k}(i)R_{k})\right]$ (16b)
@@ III-C Two-step Convex Optimization Framework
We now define the two subproblems which comprise our two-step optimization framework. It first solves Problem 4 and then uses its solution to solve Problem 5.
@@
$\hskip-85.35826ptS_{0}(i)=\rho_{0}(i){\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (18c)
where $\Sigma_{T}$ is calculated with $S_{T}(i)$ using eq. 3b
@@
Under Assumptions 1, 2 and 3, the solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$, hence, lossless.
@@
$\mathcal{L}(\cdot)=J(\Xi)+$ $\sum_{k=0}^{T-1}\left[\mathrm{Tr}(\Lambda^{\top}_{k}H_{k}^{\Sigma})+\sum_{i\in\mathcal{S}}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))\right\}\right]$ (19)
@@
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}$ $=2\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){S}_{k}^{-1}(i)=0$ (20a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}$ $=\rho_{k}(i)R_{k}-M_{k}(i)+\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (20b)
@@
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies the corresponding Lagrange multiplier $M_{k}(i)$ is singular from Lemma 1 and eq. 20c. Rearranging eq. 20a, we get $\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){S}_{k}^{-1}(i)A_{k}^{-1}(i)$. Using this and the substitutions defined earlier in eq. 10, we simplify eq. 20b as: $\rho_{k}(i)R_{k}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Taking the determinant of both sides, we get the following:
$\mathrm{det}(\rho_{k}(i)R_{k})=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))=0$
which contradicts $\rho_{k}(i)R_{k}\succ 0$ from Assumption 1. Therefore, we have $C_{k}(i)=L_{k}(i)S^{-1}_{k}(i)L_{k}^{\top}(i)-Y_{k}(i)=0,\;\forall k,i.$
@@ IV Chance-Constrained MJLS Covariance Steering
$\mathbb{P}(a^{\top}_{j}x_{k}+b_{j}\leq 0,\forall j)$ $\geq 1-\delta_{x}$ (21a)
$\mathbb{P}(\|x_{k}-\mu_{k}\|\leq d_{\max})$ $\geq 1-\epsilon_{x}$ (21b)
@@
$\mathbb{P}(f^{\top}_{j}u_{k}(i)+g_{j}\leq 0,\forall j)$ $\geq 1-\delta_{u}(i)$ (22a)
$\mathbb{P}(\|u_{k}(i)\|\leq u_{\max}(i))$ $\geq 1-\epsilon_{u}(i)$ (22b)
Previous works [23, 8, 10, 6] have focused on deterministic and convex reformulations of eqs. 21 and 22, but they usually assume state distribution to be Gaussian. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) Therefore, we utilize one-sided Chebyshev/Cantelli’s inequality[24] and the multivariate Chebyshev inequality[25] to convert the chance constraints to deterministic path constraints. Note that while the constraint on the state $x_{k}$ is mode-independent, the control chance constraints are imposed for a given mode $i$. The multi-modal nature of the control variable $u_{k}$ and practical applications motivate us to consider mode-dependent control chance constraints as it would be less conservative than mode-independent control chance constraints. For mode-conditioned control chance constraints, the mean and covariance is given by: $\mathbb{E}[u_{k}(i)|r_{k}=i]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}|r_{k}=i]=\frac{Y_{k}(i)}{\rho_{k}(i)}$ as shown in the appendix.
@@ IV-A Deterministic Formulation
Lemmas 2 and 3 are useful in the deterministic formulation of half-space and max-norm bound constraints.
Lemma 2.
Let $v\in\mathbb{R}^{n_{v}}$ be a random variable with mean $\overline{v}$ and covariance $V$. Then, $\mathbb{P}[a_{j}^{\top}{v}+b_{j}\leq 0,\forall j]\geq 1-\epsilon$ is implied by:
$a^{\top}_{j}\overline{v}+b_{j}+\sqrt{\frac{1-\epsilon_{j}}{\epsilon_{j}}a^{\top}_{j}Va_{j}}\leq 0,\quad\sum_{j}\epsilon_{j}\leq\epsilon$ (23)
Proof.
See Theorem 3.1 in [26], and Lemma 1 in [8]. ∎
Lemma 3.
Let $v\in\mathbb{R}^{n_{v}}$ be a random variable with mean $\overline{v}$ and covariance $V$. Then $\mathbb{P}[\|{v}\|\leq v_{\max}]\geq 1-\epsilon$ is implied by:
$\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}\leq v_{\max}$ (24)
Proof.
Given that $\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}\leq v_{\max}$, we know $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)})$…(A). This is described in Figure 1(a). Now using the multivariate Chebyshev inequality [25] for $v$, we get: $\mathbb{P}((v-\overline{v})^{\top}V^{-1}(v-\overline{v})\geq t^{2})\leq\frac{n_{v}}{t^{2}}$…(B). Put $t=\sqrt{\frac{n_{v}}{\epsilon}}$, and considering the complement of the event in (B), we get $\mathbb{P}((v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon})\geq 1-\epsilon$. Note that $(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}$ denotes an ellipsoid centered at $\overline{v}$ with shape matrix $V$. We first find the maximum value of $\|v\|$ inside this ellipse. Let $\mathcal{E}=\{v:(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}\}$, then: $\max_{v\in\mathcal{E}}\|v\|\leq\max_{v\in\mathcal{E}}\|v-\overline{v}\|+\|\overline{v}\|=\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}+\|\overline{v}\|$. Now construct a ball $\mathcal{B}_{v}$ centered at origin with radius $\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)}$. Geometrically, this is shown in Figure 1(b) for a $n_{v}=2$. Then, $\mathbb{P}(v\in\mathcal{B}_{v})=\mathbb{P}(\|v\|\leq\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}+\|\overline{v}\|)\geq\mathbb{P}(v\in\mathcal{E})$ since $\mathcal{E}\subseteq\mathcal{B}_{v}$. But $\mathbb{P}(v\in\mathcal{E})\geq 1-\epsilon$. Therefore, $\mathbb{P}\left(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}\right)\geq\mathbb{P}(v\in\mathcal{E})\geq 1-\epsilon$. Combining this with (A), we get $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\cdot\lambda_{\max}(V)})\geq 1-\epsilon$.
$\overline{v}||+\sqrt{\frac{n}{\epsilon}\lambda_{max}(V)}$$v_{\max}$v||v||p(){p($\cdot$)}
[…]
Figure 1: Schematics for the proof of Lemma 3. Figure (a) shows that $\mathbb{P}(\|v\|\leq v_{\max})\geq\mathbb{P}\left(\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n}{\epsilon}\lambda_{\max}(V)}\right)$. Figure (b) shows that $\mathbb{P}(v\in\mathcal{B}_{v})\geq\mathbb{P}(v\in\mathcal{E})$ where the ball $\mathcal{B}_{v}$ is given by $\{v:\|v\|\leq\|\overline{v}\|+\sqrt{\frac{n_{v}}{\epsilon}\lambda_{\max}(V)}\}$ and the ellipsoid $\mathcal{E}$ is given by $\{v:(v-\overline{v})^{\top}V^{-1}(v-\overline{v})\leq\frac{n_{v}}{\epsilon}\}$.
Remark 3.
Note that while eqs. 23 and 24 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
$\frac{1-\epsilon}{\epsilon}a^{\top}_{j}Va_{j}\leq(a_{j}^{\top}\bar{v}+b_{j})^{2}$ (25)
$\frac{n_{v}}{\epsilon}{\lambda_{\max}(V)}\leq(v_{\max}-\|\bar{v}\|)^{2}$ (26)
which are convex in the covariance $V$.